(x^2)y″+(5x)y′+4y=x^6, y(1)=8, y′(1)=0. y = ?
Answer #1

Gireen initial realue probleme is d n²y & suylt 4y = 16 . 4011=8 1111=0 net & n=e2 z=logy de dia Now, dy y & de aut 1 de laas yt uy t4y = eoz consides only the homogeneous part, - ausiliary efecatron ics - m2fumth=0 3 cm +21220 an--२ । 2 Therefore- - __ysh - +9-4 celega ___y.c17 2050 = 3 -2ch + _ 42 २८, -1- 3 ____५८ 127 fyini = ( +ps ere) + * २५ -१ 32 ।। 32 3 32- 0 08Please rate if this was helpful for you.